# IrrMap: A Large-Scale Comprehensive Dataset for Irrigation Method Mapping

### v1

@@ Abstract.
We introduce IrrMap, the first large-scale dataset (1.1 million patches) for irrigation method mapping across regions. IrrMap consists of multi-resolution satellite imagery from LandSat and Sentinel, along with key auxiliary data such as crop type, land use, and vegetation indices. The dataset spans 1,687,899 farms and 14,117,330 acres across multiple western U.S. states from 2013 to 2023, providing a rich and diverse foundation for irrigation analysis and ensuring geospatial alignment and quality control. The dataset is ML-ready, with standardized 224×224 GeoTIFF patches, the multiple input modalities, carefully chosen train-test-split data, and accompanying dataloaders for seamless deep learning model training and benchmarking in irrigation mapping. The dataset is also accompanied by a complete pipeline for dataset generation, enabling researchers to extend IrrMap to new regions for irrigation data collection or adapt it with minimal effort for other similar applications in agricultural and geospatial analysis. We also analyze the irrigation method distribution across crop groups, spatial irrigation patterns (using Shannon diversity indices), and irrigated area variations for both LandSat and Sentinel, providing insights into regional and resolution-based differences. To promote further exploration, we openly release IrrMap, along with the derived datasets, benchmark models, and pipeline code, through a GitHub repository: https://github.com/Nibir088/IrrMap and Data repository: https://huggingface.co/Nibir/IrrMap, providing comprehensive documentation and implementation details.
Remote Sensing, Irrigation Mapping, Machine Learning, Multi-modal, Multi-resolution, Satellite Imagery
ccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Creating a large-scale irrigation dataset across different regions is particularly challenging. Firstly, there is no centralized database; data for different states in the US can potentially come from different sources. Therefore, we see variations in classification systems, spatial resolutions, and labeling standards of different agencies. For instance, what is classified as “Center Pivot” irrigation in Arizona may be labeled as “Big Gun” irrigation in Washington, which requires a systematic approach to ensure consistent labeling across the region. Moreover, significant variation in farm sizes poses a challenge in balancing high-resolution mapping for small farms while ensuring comprehensive regional coverage. Textural patterns of irrigation type could be heavily influenced by crop type, soil, etc. Additional contextual information, such as crop types and vegetation indices, can be beneficial to improve classification accuracy (Dempsey, 2024; Allen et al., 2021).
@@ 1.1. Our Contributions
In this work, we introduce IrrMap, a large-scale comprehensive dataset for irrigation mapping designed to ($i$) support the development of machine learning (ML) models that identify various irrigation methods, and ($ii$) accelerate mapping various agricultural infrastructure to support resilience in food systems and water availability. Our dataset includes satellite imagery along with key variables relevant to irrigation mapping, such as crop type, land use data, and derived vegetation indices. It spans multiple states across the U.S., covering 1,668,899 farms and 11,443,492 acres from 2013 to 2023. To the best of our knowledge, IrrMap is the first and largest dataset dedicated to data-driven irrigation method mapping. As discussed below, our work addresses each of the criteria laid out in the call for paper, namely: ($i$) Accessibility: the dataset is public (Huggingface); ($ii$) Quality and Documentation: Automated and manual quality checks were conducted, along with documentation of the dataset, pipeline and analysis; ($iii$) Impact: The multi-modal dataset and derived models can be extended for national-scale mapping of irrigation and related problems as well as promote methodological advances (outlined in Section 7); ($iv$) Ethics and Fairness: Our collected data is based solely on publicly available sources and is provided for research purposes, ensuring transparency and reproducibility
Multi-resolution satellite imagery. To accommodate diverse research needs, we integrate multi-resolution data from LandSat (30m × 30m) and Sentinel (10m × 10m), ensuring a balance between detail and scalability. High-resolution imagery enables fine-grained analysis of irrigation patterns and small-scale variations, while lower-resolution data facilitates large-scale assessments with reduced computational overhead, making it ideal for regional and national studies. Besides, LandSat data is available for more years compared to Sentinel data allowing for more detailed temporal evolution studies of irrigation.
@@
Benchmarking. We also establish benchmarks using deep learning (DL) models trained on different input modalities, including RGB, RGB with Crop Type, RGB with Land Use, and RGB with NDVI, demonstrating their effectiveness in irrigation mapping tasks.
@@
Figure 2. Coverage area of the irrigation method dataset.
2. Related Work
3. Description of Data Source
@@
General Information. Our multi-resolution dataset IrrMap consists of around one million patches containing LandSat and Sentinel imagery. The dataset includes pixel-precise irrigation mapping, land use, and crop cover annotations at both 30m and 10m resolutions while preserving the geolocation of each patch (Table 1). Our dataset covers four states—Arizona, Colorado, Utah, and Washington—spanning approximately 1,443,492.31 acre of U.S. agricultural lands (shown in Figure 2 and Table 2).
@@
Irrigation Data Source. Since irrigation label data is not readily available from a single source for all states, we collected data for four states from different sources. For Utah (2021–2023), we obtained data from the Water-Related Land Use (WRLU) [^2] https://dwre-utahdnr.opendata.arcgis.com/pages/wrlu-data dataset. Washington’s irrigation data (2015–2020) was sourced from the Washington State Department of Agriculture Agricultural Land Use dataset (WSDA) [^3] https://agr.wa.gov/departments/land-and-water/natural-resources/agricultural-land-use, while Colorado’s data (2018–2020) was acquired from the Colorado Decision Support System (CDSS) [^4] https://dwr.colorado.gov/services/data-information/gis, which provides Geographic Information System (GIS) datasets for all river basins. For Arizona (2015–2017), we utilized the USGS Verified Irrigated Agricultural Lands dataset [^5] https://catalog.data.gov/dataset/verified-irrigated-agricultural-lands-for-the-united-states-200217, a GIS geodatabase developed collaboratively by USGS and the University of Wisconsin. All of these datasets include various land use details, such as vector polygons of irrigated fields, irrigation methods, crop types, water sources, and acreage information across different years. To standardize the analysis across data sources, we mapped various irrigation practices to three primary methods: drip, sprinkler, and flood irrigation (see Table 4). For example, “Center Pivot”, “Wheeler Sprinkler”, and “Traveling Gun” all fall under sprinkler irrigation. In addition, some labels are noisy as multiple irrigation methods are assigned to a single entry (e.g., “Drip/Rill/Sprinkler” or “Center Pivot/Drip/Sprinkler”). To maintain clarity, such noisy labels are removed, ensuring clean, structured data for reliable analysis. Irrigation methods vary by state; Utah’s 2.7 million acres have 0.08% drip, 34% sprinkler, and 31% flood, while Arizona’s 566,340 acres include 8% drip, 46% flood, and 20% sprinkler irrigation (See Table 1).
Crop Data Source. For Crop Data, A total of 143 distinct crop types were identified across the four states from USGS Verified Irrigated Agricultural Lands datasets. To standardize the analysis, we consolidated these crops into 20 categorical groups based on classifications from Leff et al. (Leff et al., 2004) and definition from the IR4-Project of U.S. Department of Agriculture (USDA) (Project, 2012). The groups are: Alfalfa, Cereals, Cover Crop, Fibres, Fruits, Grass, Green House, Herb Group, Horticulture, Nursery, Nuts, Oil-bearing crops, Orchard, Pulses, Roots and Tubers, Shrub Plants, Sugar Crops, Vegetables, Vineyard, and an additional category for unspecified crops. For instance, the cereal group includes barley, corn, wheat, and sorghum, while the fruit group encompasses apples, berries, citrus, and melons. The complete mapping of individual crops to their respective groups is provided in the Appendix Table 10.
@@ 4. ML-ready Dataset Preparation
Derived Indices Creation. To enhance surface property analysis for irrigation mapping, we compute various spectral indices capturing vegetation health [^6] https://www.nv5geospatialsoftware.com/docs/AlphabeticalListSpectralIndices.html, soil conditions, and water availability. NDVI, GNDVI, and CIgreen assess vegetation vigor and chlorophyll content, while EVI, SAVI, and MSAVI adjust for atmospheric and soil background effects, improving vegetation monitoring in different environments (Meivel and Maheswari, 2022; Tucker, 1985; Huete et al., 2002). NDWI detects water bodies and moisture content, aiding in drought assessment, whereas NDTI distinguishes tilled from untilled soils for land management analysis (Gao, 1996; Zheng et al., 2014). PRI measures plant stress and photosynthetic efficiency, OSAVI refines soil adjustment for vegetation detection, and WDRVI enhances differentiation in high biomass areas (Gitelson, 2004; Fern et al., 2018; Thénot et al., 2002). RVI provides an alternative vegetation measure with reduced sensitivity to atmospheric variations (Gonenc et al., 2019). Together, these indices offer valuable insights for irrigation mapping, agricultural monitoring, and environmental assessment (More details in Appendix Table 9).
@@
Train-Test Splitting. To ensure a robust and unbiased evaluation of machine learning models, we carefully split the IrriMap dataset into 80% training and 20% testing sets. As irrigation practices vary significantly vary across different geographical regions (e.g., Utah has only 0.08% drip-irrigated lands in the labeled data), a state-wise stratified split is performed. Each state’s dataset is independently divided into an 80-20% split, and the resulting 20% testing portions are then combined to form a spatially well-represented test set, thereby mitigating geographic biases. State-wise distribution of training and testing datasets is shown in Table 3. A total of 934,514 (88,015 for LandSat and 846,499 for Sentinel) patches are created for the training dataset. The test dataset has 233,645 (22,011 for LandSat and 211,634 for Sentinel) patches in total.
@@ 6. Benchmarking
Table 5. Irrigation Classification Performance for Different Modalities, reporting Precision (P), Recall (R), F1-score (F), and Intersection over Union (IoU) for Flood, Sprinkler, and Drip irrigation types across different states. More results on Sentinel data are in Appendix D Table 8.
State Modality Flood Sprinkler Drip
@@
To evaluate the effectiveness of IrrMap for irrigation method classification, we designed an experiment using satellite imagery from Landsat. Our goal is to classify irrigation types—Flood, Sprinkler, and Drip—by leveraging different input modalities and benchmarking their performance across various configurations. We employed a UNet-style ResNet-34 architecture for all baseline experiments, training the models on the complete Landsat training dataset and presenting results on the test set of each state. We train the model for 10 epochs on an NVIDIA A100 80GB GPU. The motivation behind this experiment is to assess how incorporating auxiliary modalities—such as NDVI, Crop Mask, and Land Mask—alongside RGB imagery improves impact the performance. More hyper-parameter settings and results on the Sentinel data are present in Appendix D Table 8.
@@
The results demonstrate that incorporating additional feature layers alongside RGB imagery leads to varying degrees of improvement in irrigation classification performance. RGB + Crop Mask consistently achieves the highest accuracy, with an average F1-score improvement of 30–50% over RGB alone, particularly excelling in Flood and Sprinkler irrigation. The Land Mask augmentation also enhances classification, yielding a 20–40% increase in IoU and F1-score compared to RGB-only, especially for Flood irrigation. Meanwhile, NDVI augmentation provides moderate benefits, with a 10–25% boost in F1-score and IoU for Sprinkler and Flood irrigation methods but does not perform as well for Drip irrigation. RGB-only models still produce reasonable results, particularly for Flood irrigation, but exhibit relatively lower performance in distinguishing Sprinkler and Drip irrigation methods. These findings suggest that crop-specific information (via Crop Mask) is the most informative additional layer, while land use and vegetation indices (NDVI) contribute moderately, depending on the irrigation methods. This opens up an interesting research direction on how to best leverage these additional modalities—whether through feature fusion techniques, advanced model architectures, or task-specific weighting—to achieve optimal performance.

### v2

@@ Abstract.
We introduce IrrMap, the first large-scale dataset (1.1 million patches) for irrigation method mapping across regions. IrrMap consists of multi-resolution satellite imagery from LandSat and Sentinel, along with key auxiliary data such as crop type, land use, and vegetation indices. The dataset spans 1,687,899 farms and 14,117,330 acres across multiple western U.S. states from 2013 to 2023, providing a rich and diverse foundation for irrigation analysis and ensuring geospatial alignment and quality control. The dataset is ML-ready, with standardized 224×224 GeoTIFF patches, the multiple input data layers, carefully chosen train-test-split data, and accompanying dataloaders for seamless deep learning model training and benchmarking in irrigation mapping. The dataset is also accompanied by a complete pipeline for dataset generation, enabling researchers to extend IrrMap to new regions for irrigation data collection or adapt it with minimal effort for other similar applications in agricultural and geospatial analysis. We also analyze the irrigation method distribution across crop groups, spatial irrigation patterns (using Shannon diversity indices), and irrigated area variations for both LandSat and Sentinel, providing insights into regional and resolution-based differences. To promote further exploration, we openly release IrrMap, along with the derived datasets, benchmark models, and pipeline code, through a GitHub repository: https://github.com/Nibir088/IrrMap and Data repository: https://huggingface.co/Nibir/IrrMap, providing comprehensive documentation and implementation details.
Remote Sensing; Irrigation Mapping; Machine Learning; Multi-source; Multi-resolution; Satellite Imagery
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737380isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Machine learningccs: Computing methodologies Learning paradigmsccs: Computing methodologies Computer visionccs: Applied computing Agriculture
1. Introduction
Figure 1. Spatial diversity and structural heterogeneity of irrigation: Multiple irrigation types can co-occur in small areas along with non-irrigated areas. Even among fields with the same type, we see diversity in the way irrigation is implemented (for e.g., central pivot vs. non-central pivot). The spatial resolution of the imagery sources can adequately represent small fields as well (see 2nd row).
@@ 1. Introduction
Creating a large-scale irrigation dataset across different regions is particularly challenging. Firstly, there is no centralized database; data for different states in the US can potentially come from different sources. Therefore, we see variations in classification systems, spatial resolutions, and labeling standards of different agencies (See Figure 1). For instance, what is classified as “Center Pivot” irrigation in Arizona may be labeled as “Big Gun” irrigation in Washington, which requires a systematic approach to ensure consistent labeling across the region. Moreover, significant variation in farm sizes poses a challenge in balancing high-resolution mapping for small farms while ensuring comprehensive regional coverage. Textural patterns of irrigation type could be heavily influenced by crop type, soil, etc. Additional contextual information, such as crop types and vegetation indices, can be beneficial to improve classification accuracy (Dempsey, 2024; Allen et al., 2021).
@@ 1.1. Our Contributions
In this work, we introduce IrrMap, a large-scale comprehensive dataset for irrigation mapping designed to ($i$) support the development of machine learning (ML) models that identify various irrigation methods, and ($ii$) accelerate mapping various agricultural infrastructure to support resilience in food systems and water availability. Our dataset includes satellite imagery along with key variables relevant to irrigation mapping, such as crop type, land use data, and derived vegetation indices. It spans multiple states across the U.S., covering 1,668,899 farms and 11,443,492 acres from 2013 to 2023. To the best of our knowledge, IrrMap is the first and largest dataset dedicated to data-driven irrigation method mapping. As discussed below, our work addresses each of the criteria laid out in the call for paper, namely: ($i$) Accessibility: the dataset is public (Huggingface); ($ii$) Quality and Documentation: Automated and manual quality checks were conducted, along with documentation of the dataset, pipeline and analysis; ($iii$) Impact: The multi-source dataset and derived models can be extended for national-scale mapping of irrigation and related problems as well as promote methodological advances (outlined in Section 7); ($iv$) Ethics and Fairness: Our collected data is based solely on publicly available sources and is provided for research purposes, ensuring transparency and reproducibility
Multi-resolution satellite imagery. To accommodate diverse research needs, we integrate multi-resolution data from LandSat (30m × 30m) and Sentinel (10m × 10m), ensuring a balance between detail and scalability. High-resolution imagery enables fine-grained analysis of irrigation patterns and small-scale variations, while lower-resolution data facilitates large-scale assessments with reduced computational overhead, making it ideal for regional and national studies. Besides, LandSat data is available for more years compared to Sentinel data allowing for more detailed temporal evolution studies of irrigation. While LandSat and Sentinel are at coarser spatial resolution, the spatial and temporal coverages of these sources is much higher than the publicly avaliable higher resolution images.
@@
Benchmarking. We also establish benchmarks using deep learning (DL) models trained on different input layers, including RGB, RGB with Crop Type, RGB with Land Use, and RGB with NDVI, demonstrating their effectiveness in irrigation mapping tasks.
@@
Figure 3. Spatial and temporal distribution of irrigation dataset. (a) Statistical summary of irrigation type coverage by state, showing percentage distribution across irrigation types, total irrigated area, and farm count. Note the significant class imbalance across irrigation categories. (b) Geographic distribution of study areas within the United States.
@@
2. Related Work
3. Description of Data Source
General Information. Our multi-resolution dataset IrrMap consists of around one million patches containing LandSat and Sentinel imagery. The dataset includes pixel-precise irrigation mapping, land use, and crop cover annotations at both 30m and 10m resolutions while preserving the geolocation of each patch (Table 3(a)). Our dataset covers four states—Arizona, Colorado, Utah, and Washington—spanning approximately 1,443,492.31 acre of U.S. agricultural lands (shown in Figure 3(b) and Table 1).
@@ 3. Description of Data Source
Irrigation Data Source. Since irrigation label data is not readily available from a single source for all states, we collected data for four states from different sources. For Utah (2021–2023), we obtained data from the Water-Related Land Use (WRLU) [^2] https://dwre-utahdnr.opendata.arcgis.com/pages/wrlu-data dataset. Washington’s irrigation data (2015–2020) was sourced from the Washington State Department of Agriculture Agricultural Land Use dataset (WSDA) [^3] https://agr.wa.gov/departments/land-and-water/natural-resources/agricultural-land-use, while Colorado’s data (2018–2020) was acquired from the Colorado Decision Support System (CDSS) [^4] https://dwr.colorado.gov/services/data-information/gis, which provides Geographic Information System (GIS) datasets for all river basins. For Arizona (2015–2017), we utilized the USGS Verified Irrigated Agricultural Lands dataset [^5] https://catalog.data.gov/dataset/verified-irrigated-agricultural-lands-for-the-united-states-200217, a GIS geodatabase developed collaboratively by USGS and the University of Wisconsin. All of these datasets include various land use details, such as vector polygons of irrigated fields, irrigation methods, crop types, water sources, and acreage information across different years. To standardize the analysis across data sources, we mapped various irrigation practices to three primary methods: drip, sprinkler, and flood irrigation (see Table 3). For example, “Center Pivot”, “Wheeler Sprinkler”, and “Traveling Gun” all fall under sprinkler irrigation. In addition, some labels are noisy as multiple irrigation methods are assigned to a single entry (e.g., “Drip/Rill/Sprinkler” or “Center Pivot/Drip/Sprinkler”). To maintain clarity, such noisy labels are removed, ensuring clean, structured data for reliable analysis. Irrigation methods vary by state; Utah’s 2.7 million acres have 0.08% drip, 34% sprinkler, and 31% flood, while Arizona’s 566,340 acres include 8% drip, 46% flood, and 20% sprinkler irrigation (See Table 3(a)).
Crop Data Source. For Crop Data, A total of 143 distinct crop types were identified across the four states from USGS Verified Irrigated Agricultural Lands datasets. To standardize the analysis, we consolidated these crops into 20 categorical groups based on classifications from Leff et al. (Leff et al., 2004) and definition from the IR4-Project of U.S. Department of Agriculture (USDA) (Project, 2012). The groups are: Alfalfa, Cereals, Cover Crop, Fibres, Fruits, Grass, Green House, Herb Group, Horticulture, Nursery, Nuts, Oil-bearing crops, Orchard, Pulses, Roots and Tubers, Shrub Plants, Sugar Crops, Vegetables, Vineyard, and an additional category for unspecified crops. For instance, the cereal group includes barley, corn, wheat, and sorghum, while the fruit group encompasses apples, berries, citrus, and melons. The complete mapping of individual crops to their respective groups is provided in the Appendix.
@@ 4. ML-ready Dataset Preparation
Derived Indices Creation. To enhance surface property analysis for irrigation mapping, we compute various spectral indices capturing vegetation health [^6] https://www.nv5geospatialsoftware.com/docs/AlphabeticalListSpectralIndices.html, soil conditions, and water availability. NDVI, GNDVI, and CIgreen assess vegetation vigor and chlorophyll content, while EVI, SAVI, and MSAVI adjust for atmospheric and soil background effects, improving vegetation monitoring in different environments (Meivel and Maheswari, 2022; Tucker, 1985; Huete et al., 2002). NDWI detects water bodies and moisture content, aiding in drought assessment, whereas NDTI distinguishes tilled from untilled soils for land management analysis (Gao, 1996; Zheng et al., 2014). PRI measures plant stress and photosynthetic efficiency, OSAVI refines soil adjustment for vegetation detection, and WDRVI enhances differentiation in high biomass areas (Gitelson, 2004; Fern et al., 2018; Thénot et al., 2002). RVI provides an alternative vegetation measure with reduced sensitivity to atmospheric variations (Gonenc et al., 2019). Together, these indices offer valuable insights for irrigation mapping, agricultural monitoring, and environmental assessment (More details in Appendix.
@@
Train-Test Splitting. To ensure a robust and unbiased evaluation of machine learning models, we carefully split the IrriMap dataset into 80% training and 20% testing sets. As irrigation practices vary significantly vary across different geographical regions (e.g., Utah has only 0.08% drip-irrigated lands in the labeled data), a state-wise stratified split is performed. Each state’s dataset is independently divided into an 80-20% split, and the resulting 20% testing portions are then combined to form a spatially well-represented test set, thereby mitigating geographic biases. State-wise distribution of training and testing datasets is shown in Table 2. A total of 934,514 (88,015 for LandSat and 846,499 for Sentinel) patches are created for the training dataset. The test dataset has 233,645 (22,011 for LandSat and 211,634 for Sentinel) patches in total. To improve accessibility for downstream users, we additionally provide a lightweight version of the training set, containing 234,861 Sentinel and 26,316 LandSat patches. These subsets were created by randomly sampling 30% of the unique satellite scene tiles and retaining all associated patches.
Figure 4. Our designed user interface for quality filtering at the patch level.
@@ 6. Benchmarking
Table 4. Irrigation Classification Performance for Different Input Layers, reporting Precision (P), Recall (R), F1-score (F), and Intersection over Union (IoU) for Flood, Sprinkler, and Drip irrigation types across different states.
State Input Layers Flood Sprinkler Drip
@@
Table 5. Irrigation Classification Performance for Different Input layers, reporting Precision (P), Recall (R), F1-score (F), and Intersection over Union (IoU) for Flood, Sprinkler, and Drip irrigation methods across LandSat and Sentinel data.
Satellite Input Layers Flood Sprinkler Drip
P R F IoU P R F IoU P R F IoU
Evaluation Setup. To evaluate the effectiveness of IrrMap for irrigation method classification, we design an experiment using satellite imagery from LandSat. In particular, we demonstrate how additional masks can contribute to performance improvements as well as quantify the impact of each of them. We select three primary layers (CROP, LAND and NDVI) and explore the extent to which each one enhanced the results.
@@
We classify irrigation types—Flood, Sprinkler, and Drip—by leveraging different input layers and benchmarking their performance across various configurations. We employed a UNet-style ResNet-34 architecture for all baseline experiments, training the models on the complete LandSat training dataset and presenting results on the test set of each state. We train the model for 10 epochs on an NVIDIA A100 80GB GPU. The motivation behind this experiment is to assess how incorporating auxiliary layers—such as NDVI, Crop Mask, and Land Mask—alongside RGB imagery improves impact the performance. More hyper-parameter settings and results on the Sentinel data are present in Appendix.
@@
Results. The results in Table 4 demonstrate that incorporating additional feature layers alongside RGB imagery leads to varying degrees of improvement in irrigation classification performance. RGB + Crop Mask consistently achieves the highest accuracy, with an average F1-score improvement of 30–50% over RGB alone, particularly excelling in Flood and Sprinkler irrigation. The Land Mask augmentation also enhances classification, yielding a 20–40% increase in IoU and F1-score compared to RGB-only, especially for Flood irrigation. Meanwhile, NDVI augmentation provides moderate benefits, with a 10–25% boost in F1-score and IoU for Sprinkler and Flood irrigation methods but does not perform as well for Drip irrigation. RGB-only models still produce reasonable results, particularly for Flood irrigation, but exhibit relatively lower performance in distinguishing Sprinkler and Drip irrigation methods. These findings suggest that crop-specific information (via Crop Mask) is the most informative additional layer, while land use and vegetation indices (NDVI) contribute moderately, depending on the irrigation methods. This opens up an interesting research direction on how to best leverage these additional layers—whether through feature fusion techniques, advanced model architectures, or task-specific weighting—to achieve optimal performance.
In Table 5, we present the overall performance on the training set. For training on Sentinel data, we use the same hyper-parameters and model architecture as LandSat. We randomly sample 100K training and test data from the IrrMap Sentinel dataset and train the model for five epochs.
The results indicate that the inclusion of additional bands, such as crop type and NDVI, has a notable impact on irrigation classification performance for both LandSat and Sentinel datasets. In LandSat, the RGB+CROP layers consistently outperforms other configurations, particularly in Flood and Sprinkler irrigation, suggesting that crop type information significantly enhances classification accuracy. Similarly, RGB+NDVI improves performance compared to standard RGB, but it does not surpass RGB+CROP, highlighting the greater influence of crop data over vegetation indices in LandSat imagery. For Sentinel, the impact of additional bands is even more pronounced, with RGB+CROP achieving the highest F1-scores and IoU across all irrigation methods, particularly for Sprinkler and Drip irrigation, where detailed crop information enhances model differentiation. RGB+LAND has also shown similar impact. The RGB+NDVI configuration also improves performance, especially in Sprinkler irrigation, reinforcing the value of vegetation indices at Sentinel’s higher resolution. Overall, the inclusion of auxiliary bands positively influences classification performance, with crop type data playing a critical role in both datasets.